OpenStudy (anonymous):

If f(x) = 3x – 1 and g(x) = x2 – 2x – 2, what is f(g(x))

OpenStudy (anonymous):

F(x) = 3x + 1 g(x) = x^2 + 2x + 1 f(g(x)) = 3(x^2 + 2x + 1) + 1 = 3x^2 + 6x + 4

OpenStudy (anonymous):

can i get a woot woot! - <3 seashell

OpenStudy (anonymous):

woot woot

OpenStudy (anonymous):

;D yaaaaaay!!:D